OpenStudy (anonymous):

Write in the Complex Form (a+bi) (5+i)^2

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

@bohotness

OpenStudy (mathstudent55):

First, square the binomial. Either use the square of a binomial formula, or use FOIL. Then collect like terms and write it in a + bi form.

OpenStudy (mathstudent55):

Here is the formula you need. \((a + b)^2 = a^2 + 2ab + b^2\)

OpenStudy (anonymous):

i got 25+10i+i^2 but idk how to write that in complex form

OpenStudy (anonymous):

@mathstudent55

ganeshie8 (ganeshie8):

Hint : \[\large i^2 = -1\]

OpenStudy (anonymous):

24+10i

ganeshie8 (ganeshie8):

thats it!

OpenStudy (anonymous):

could i ask you one more question?

ganeshie8 (ganeshie8):

sure ask

OpenStudy (anonymous):

i-6 --- i

ganeshie8 (ganeshie8):

First, divide \(i\) by each term in the numerator

OpenStudy (anonymous):

i thought you would multiply it

ganeshie8 (ganeshie8):

\[\large \dfrac{i-6}{i} = \dfrac{i}{i} - \dfrac{6}{i} = 1 - \dfrac{6}{i}\]

ganeshie8 (ganeshie8):

you're right, now multiply \(i\) top and bottom of the second term

ganeshie8 (ganeshie8):

\[\large 1 - \dfrac{6\color{gray}{\times i}}{i\color{gray}{\times i}}\]

ganeshie8 (ganeshie8):

\[\large 1 - \dfrac{6i}{i^2}\]

ganeshie8 (ganeshie8):

\[\large 1 - \dfrac{6i}{-1}\]

ganeshie8 (ganeshie8):

\[\large 1 +6i\]

OpenStudy (anonymous):

thank you!

ganeshie8 (ganeshie8):

just so you know you may multiply the conjugate top and bottom of the given expression directly as well..